# Dynamics of the general $Q$-tensor model interacting with a rigid body

### v1

@@ 1. Introduction
In this work, we investigate the fluid-rigid body interaction problem of the general Beris-Edwards $Q$-tensor model [7]. More precisely, we show the local strong well-posedness for large data and the global strong well-posedness for small data of the interaction problem. In particular, in the underlying $Q$-tensor model, we consider an arbitrary ratio of tumbling and alignment effects $\xi\in\mathbb{R}$. For the fluid velocity $u$ and pressure $p$ and the molecular orientation of the liquid crystals $Q$, taking values in the space $\mathbb{S}_{0}^{3}$ of symmetric and traceless $3\times 3$-matrices, the general $Q$-tensor model is of the form
[…]
@@
$\tau(Q,H(Q))$ $=-\nabla Q\odot\nabla Q-\xi(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})H(Q)-\xi H(Q)(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})+2\xi(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}\bigl{(}QH(Q)\bigr{)}\enspace\text{and}\enspace$
@@
(1.2) $m_{\mathcal{S}}h^{\prime\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}\Sigma(u,\pi,Q)\nu(t)\,\mathrm{d}\Gamma\enspace\text{and}\enspace(J\Omega)^{\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}(x-h(t))\times\Sigma(u,\pi,Q)\nu(t)\,\mathrm{d}\Gamma.$
@@
(1.3) $u(t,x)=h^{\prime}(t)+\Omega(t)\times(x-h(t)),\enspace\text{for}\enspace t\in(0,T),\enspace x\in\partial\mathcal{S}(t).$
@@
We establish the maximal $\mathrm{L}^{p}$-regularity on the anisotropic space $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{3}\times\mathbb{R}^{3}$, where $\mathcal{F}_{0}$ is the initial fluid domain, and $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ denotes the space of solenoidal vector fields in $\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$. In particular, the condition $Q\in\mathbb{S}_{0}^{3}$ is included in the function spaces in an elegant way. The maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity with exponential decay for the linearization at zero follows from spectral theory.
In order to obtain the first main result of this paper, the local strong well-posedness of the interaction problem (1.1)–(1.3) for large initial data made precise below in Theorem 4.1, we estimate the nonlinear terms and employ the maximal regularity. The anisotropic ground space $\mathrm{L}^{2}\times\mathrm{H}^{1}$ for the liquid crystals part renders the nonlinear problem more complicated.
With regard to the second main result of the present article, the global strong well-posedness and exponential decay of the solution for small initial data as asserted in Theorem 4.2, we first prove that the total energy of the interaction problem (1.1)–(1.3) is decreasing, which reveals the shape of the equilibria. In a second step, we make use of a fixed point argument, relying on the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity of the linearization at zero as well as adjusted decay estimates of the transform.
@@ 2. The fluid-rigid body interaction problem of the general $Q$-tensor model
First, we describe the general $Q$-tensor model. By $\mathcal{F}_{T}$, we denote the spacetime associated to the fluid domain and made precise below. The variables are the fluid velocity $u^{\mathcal{F}}\colon\mathcal{F}_{T}\to\mathbb{R}^{3}$, the pressure $\pi^{\mathcal{F}}\colon\mathcal{F}_{T}\to\mathbb{R}$ and the molecular orientation of the liquid crystal $Q^{\mathcal{F}}\colon\mathcal{F}_{T}\to\mathbb{S}_{0,\mathbb{R}}^{3}$. Here $\mathbb{S}_{0,\mathbb{R}}^{3}$ denotes the space of symmetric $3\times 3$-matrices with trace zero, so $\mathbb{S}_{0,\mathbb{R}}^{3}\coloneqq\{M\in\mathbb{R}^{3\times 3}:M=M^{\top}\enspace\text{and}\enspace\operatorname{tr}M=0\}$. The complex version $\mathbb{S}_{0,\mathbb{C}}^{3}$ is defined analogously. We also write $\mathbb{S}_{0}^{3}$ if no confusion arises. For $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$, we denote by $\mathbb{M}_{0,\mathbb{K}}^{3}\coloneqq\{M\in\mathbb{K}^{3\times 3}:\operatorname{tr}M=0\}$ the space of traceless matrices in $\mathbb{K}^{3\times 3}$ and also occasionally omit the subscript K. More background on $Q$-tensor models, including the relation with the Ericksen-Leslie model, can be found in [6], see also [21, Section 2].
One feature of the $Q$-tensor model is the free energy functional taking the shape
@@
$\quad+2\xi(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}\bigl{(}Q^{\mathcal{F}}H(Q^{\mathcal{F}})\bigr{)}\enspace\text{and}\enspace$
@@
For simplicity, we assume the parameters to be equal to $1$ in the sequel, so $\mu=\lambda=a=b=c=1$, because it does not affect the analysis. The according Beris-Edwards model is given by
(2.5) ${\partial_{t}u^{\mathcal{F}}}+\left(u^{\mathcal{F}}\cdot\nabla\right)u^{\mathcal{F}}+\nabla\pi^{\mathcal{F}}&=\Delta u^{\mathcal{F}}+\operatorname{div}\bigl{(}\tau(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})\bigr{)},\enspace\operatorname{div}u^{\mathcal{F}}=0,&&\enspace\text{in}\enspace\mathcal{F}_{T},\\ \partial_{t}Q^{\mathcal{F}}+\left(u^{\mathcal{F}}\cdot\nabla\right)Q^{\mathcal{F}}-S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})&=H(Q^{\mathcal{F}}),&&\enspace\text{in}\enspace\mathcal{F}_{T}.\end{aligned}\right.$
@@
$u^{\mathcal{S}}(t,x)=h^{\prime}(t)+\Omega(t)\times(x-h(t)),\enspace\text{for all}\enspace(t,x)\in\mathcal{S}_{T},$
@@
$J_{0}=\int\limits_{\mathcal{S}_{0}}\left(|x|^{2}\operatorname{Id}_{3}-x\otimes x\right)\,\mathrm{d}x\enspace\text{as well as}\enspace J(t)=\int\limits_{\mathcal{S}(t)}\left(|x-h(t)|^{2}\operatorname{Id}_{3}-(x-h(t))\otimes(x-h(t))\right)\,\mathrm{d}x.$
@@
(2.6) $J(t)=\mathbb{O}(t)J_{0}\mathbb{O}^{\top}(t)\enspace\text{and}\enspace J(t)x_{1}\cdot x_{2}=\int\limits_{\mathcal{S}_{0}}(x_{1}\times\mathbb{O}(t)y)\cdot(x_{2}\times\mathbb{O}(t)y)\,\mathrm{d}y,\enspace\text{for all}\enspace x_{1},x_{2}\in\mathbb{R}^{3}.$
With the stress tensor $\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})$ from (2.2), the motion of the rigid body follows Newton’s law, so
(2.7) $m_{\mathcal{S}}h^{\prime\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})\nu(t)\,\mathrm{d}\Gamma,\enspace(J\Omega)^{\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}(x-h(t))\times\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})\nu(t)\,\mathrm{d}\Gamma,$
where $\nu(t)$ represents the unit outward normal to the boundary of $\mathcal{F}(t)$, so it is directed towards $\mathcal{S}(t)$.
@@
(2.8) $u^{\mathcal{F}}=0,\enspace\partial_{\nu}Q^{\mathcal{F}}=0,\enspace\text{on}\enspace(0,T)\times\partial\mathcal{O},\enspace\text{and}\enspace u^{\mathcal{F}}=u^{\mathcal{S}},\enspace\partial_{\nu}Q^{\mathcal{F}}=0,\enspace\text{on}\enspace\partial\mathcal{S}_{T}.$
@@
(2.9) $u^{\mathcal{F}}(0,x)=v_{0}(x),\enspace Q^{\mathcal{F}}(0,x)=Q_{0}(x),\enspace\text{in}\enspace\mathcal{F}_{0},\enspace h(0)=0,\enspace h^{\prime}(0)=\ell_{0}\enspace\text{and}\enspace\Omega(0)=\omega_{0}.$
@@
In this paper, $\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ represents the $\mathrm{L}^{2}(\mathcal{F}_{0})$-functions with spatial average zero, while $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ denotes the solenoidal vector fields on $\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$, i. e., for the Helmholtz projection $\mathbb{P}$, we define $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\coloneqq\mathbb{P}\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$. For $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})$, $\mathrm{Y}_{\gamma}\coloneqq\mathrm{B}_{2p}^{2-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0})^{3}\cap\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{B}_{2p}^{3-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{3}\times\mathbb{R}^{3}$ and $p>4$, we then set
(2.10) $\mathrm{X}_{\gamma}\coloneqq\{z_{0}\in\mathrm{Y}_{\gamma}:v_{0}=0,\enspace\partial_{\nu}Q_{0}=0,\enspace\text{on}\enspace\partial\mathcal{O},\enspace v_{0}=\ell_{0}+\omega_{0}\times y,\enspace\partial_{\nu}Q_{0}=0,\enspace\text{on}\enspace\partial\mathcal{S}_{0}\},$
where $\mathrm{B}_{2p}^{s}(\mathcal{F}_{0})$ denotes the Besov space. The traces and normal derivatives are well-defined due to $p>4$.

### v2

@@ 1. Introduction
In this work, we investigate the fluid-rigid body interaction problem of the general Beris-Edwards $Q$-tensor model [7]. More precisely, we show the local strong well-posedness for large data and the global strong well-posedness for small data of the interaction problem. In particular, in the underlying $Q$-tensor model, we consider an arbitrary ratio of tumbling and alignment effects $\xi\in\mathbb{R}$. For the fluid velocity $u$ and pressure $\pi$ and the molecular orientation of the liquid crystals $Q$, taking values in the space $\mathbb{S}_{0}^{3}$ of symmetric and traceless $3\times 3$-matrices, the general $Q$-tensor model is of the form
(1.1) $\left\{\begin{aligned} \partial_{t}u+\left(u\cdot\nabla\right)u+\nabla\pi&=\Delta u+\operatorname{div}\bigl(\tau(Q,H(Q))+\sigma(Q,H(Q))\bigr),\kern 5.0pt\operatorname{div}u=0,\\ \partial_{t}Q+\left(u\cdot\nabla\right)Q-S(\nabla u,Q)&=H(Q).\end{aligned}\right$.
@@
$\tau(Q,H(Q))$ $=-\nabla Q\odot\nabla Q-\xi(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})H(Q)-\xi H(Q)(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})+2\xi(Q+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}\bigl(QH(Q)\bigr)\kern 5.0pt\text{and}\kern 5.0pt$
@@
(1.2) $m_{\mathcal{S}}h^{\prime\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}\Sigma(u,\pi,Q)\nu(t)\,\mathrm{d}\Gamma\kern 5.0pt\text{and}\kern 5.0pt(J\Omega)^{\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}(x-h(t))\times\Sigma(u,\pi,Q)\nu(t)\,\mathrm{d}\Gamma.$
@@
(1.3) $u(t,x)=h^{\prime}(t)+\Omega(t)\times(x-h(t)),\kern 5.0pt\text{for}\kern 5.0ptt\in(0,T),\kern 5.0ptx\in\partial\mathcal{S}(t).$
@@
We establish the maximal $\mathrm{L}^{p}$-regularity on the anisotropic space $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{H}^{1}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{3}\times\mathbb{R}^{3}$, where $\mathcal{F}_{0}$ is the initial fluid domain, and $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ denotes the space of solenoidal vector fields in $\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$. In particular, the condition $Q\in\mathbb{S}_{0}^{3}$ is included in the function spaces in elegant way. The maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity with exponential decay for the linearization at zero follows from spectral theory.
In order to obtain the first main result of this paper, the local strong well-posedness of the interaction problem (1.1)–(1.3) for large initial data made precise below in Theorem 4.1, we estimate the nonlinear terms and employ the maximal regularity. The anisotropic ground space $\mathrm{L}^{2}\times\mathrm{H}^{1}$ for the liquid crystals part renders the nonlinear problem more complicated. The time interval of existence in this case is limited by the requirement that we exclude the case of collision of the rigid body with the outer fluid boundary.
With regard to the second main result of the present article, the global strong well-posedness and exponential decay of the solution for small initial data as asserted in Theorem 4.2, we first prove that the total energy of the interaction problem (1.1)–(1.3) is decreasing, which reveals the shape of the equilibria. In a second step, we make use of a fixed point argument, relying on the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity of the linearization at zero as well as adjusted decay estimates of the transform. Let us emphasize that we are especially able to show that no collision of the rigid body with the outer fluid boundary occurs. This follows from the exponential stability of the semigroup associated to the linearized problem around zero and the smallness of the initial data. The exponential stability also reveals that all velocities tend to zero as time tends to infinity. However, it remains an interesting open problem to determine if the position of the center of the moving rigid body converges to some point $h_{\infty}\in\mathbb{R}^{3}$ as $t\rightarrow\infty$. In the situation of a rigid body immersed in a viscous, incompressible Newtonian fluid filling the entire $\mathbb{R}^{3}$, this problem has recently been tackled in the case of a ball in [15], then for a rigid body of arbitrary shape in [28] and for the case of several bodies in [10].
@@
For convenience of the reader, let us give a brief, non-technical overview of the procedure to obtain the well-posedness results of the present interaction problem of a rigid body with the general Beris-Edwards $Q$-tensor model. First, we determine the set of equilibrium solutions by examining the critical points of the energy functional associated with the interaction problem. Next, we employ a transform that only acts locally, i. e., in a neighborhood of the translating and rotating rigid body. By continuity of the body velocities, at least for short time, we can guarantee that there is no collision of the rigid body with the outer fluid boundary. After transforming the moving boundary problem to a fixed domain, we perform the analysis of the linearized problem. The key idea here is to use a lifting procedure in order to eliminate the pressure from the surface integrals of the stress tensor, and to reformulate the linearized problem in operator form. This is done by means of the so-called added mass operator. The idea of this concept can be traced back to Bessel, see [41, 42], and has become classical in the investigation of fluid-rigid body interaction problems. Thanks to a refined spectral analysis, we prove the exponential stability for the linearization at zero, which, together with the smallness of the initial data, helps to show that no collision of the rigid body with the outer fluid boundary occurs for sufficiently small initial data. The proofs of the well-posedness results are based on a fixed point procedure, relying in turn on the analysis of the linearized problem, and they are complemented by suitable nonlinear estimates, where we efficiently handle the transform. Let us emphasize that the complicated nature of the stress tensor, in particular in the terms with the molecular orientation $Q$, leads to involved terms that need to be estimated.
@@ 2. The fluid-rigid body interaction problem of the general $Q$-tensor model
First, we describe the general $Q$-tensor model. By $\mathcal{F}_{T}$, we denote the spacetime associated to the fluid domain and made precise below. The variables are the fluid velocity $u^{\mathcal{F}}\colon\mathcal{F}_{T}\to\mathbb{R}^{3}$, the pressure $\pi^{\mathcal{F}}\colon\mathcal{F}_{T}\to\mathbb{R}$ and the molecular orientation of the liquid crystal $Q^{\mathcal{F}}\colon\mathcal{F}_{T}\to\mathbb{S}_{0,\mathbb{R}}^{3}$. Here $\mathbb{S}_{0,\mathbb{R}}^{3}$ denotes the space of symmetric $3\times 3$-matrices with trace zero, so $\mathbb{S}_{0,\mathbb{R}}^{3}\coloneqq\{M\in\mathbb{R}^{3\times 3}:M=M^{\top}\kern 5.0pt\text{and}\kern 5.0pt\operatorname{tr}M=0\}$. The complex version $\mathbb{S}_{0,\mathbb{C}}^{3}$ is defined analogously. We also write $\mathbb{S}_{0}^{3}$ if no confusion arises. For $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$, we denote by $\mathbb{M}_{0,\mathbb{K}}^{3}\coloneqq\{M\in\mathbb{K}^{3\times 3}:\operatorname{tr}M=0\}$ the space of traceless matrices in $\mathbb{K}^{3\times 3}$ and also occasionally omit the subscript K. More background on $Q$-tensor models, including the relation with the Ericksen-Leslie model, can be found in [6], see also [22, Section 2].
One feature of the $Q$-tensor model is the free energy functional taking the form
@@
$\quad+2\xi(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}\bigl(Q^{\mathcal{F}}H(Q^{\mathcal{F}})\bigr)\kern 5.0pt\text{and}\kern 5.0pt$
@@
For simplicity, we assume the parameters to be equal to $1$ in the sequel, so $\mu=\lambda=a=b=c=1$, because it does not affect the analysis. The resulting Beris-Edwards model is given by
(2.5) $\left\{\begin{aligned} {\partial_{t}u^{\mathcal{F}}}+\left(u^{\mathcal{F}}\cdot\nabla\right)u^{\mathcal{F}}+\nabla\pi^{\mathcal{F}}&=\Delta u^{\mathcal{F}}+\operatorname{div}\bigl(\tau(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})\bigr),\kern 5.0pt\operatorname{div}u^{\mathcal{F}}=0,&&\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{T},\\ \partial_{t}Q^{\mathcal{F}}+\left(u^{\mathcal{F}}\cdot\nabla\right)Q^{\mathcal{F}}-S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})&=H(Q^{\mathcal{F}}),&&\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{T}.\end{aligned}\right$.
@@
$u^{\mathcal{S}}(t,x)=h^{\prime}(t)+\Omega(t)\times(x-h(t)),\kern 5.0pt\text{for all}\kern 5.0pt(t,x)\in\mathcal{S}_{T},$
@@
$J_{0}=\int\limits_{\mathcal{S}_{0}}\left(|x|^{2}\operatorname{Id}_{3}-x\otimes x\right)\,\mathrm{d}x\kern 5.0pt\text{as well as}\kern 5.0ptJ(t)=\int\limits_{\mathcal{S}(t)}\left(|x-h(t)|^{2}\operatorname{Id}_{3}-(x-h(t))\otimes(x-h(t))\right)\,\mathrm{d}x.$
@@
(2.6) $J(t)=\mathbb{O}(t)J_{0}\mathbb{O}^{\top}(t)\kern 5.0pt\text{and}\kern 5.0ptJ(t)x_{1}\cdot x_{2}=\int\limits_{\mathcal{S}_{0}}(x_{1}\times\mathbb{O}(t)y)\cdot(x_{2}\times\mathbb{O}(t)y)\,\mathrm{d}y,\kern 5.0pt\text{for all}\kern 5.0ptx_{1},x_{2}\in\mathbb{R}^{3}.$
Using the stress tensor $\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})$ from (2.2), and noting that $\nu(t)$ denotes the unit outward normal to the boundary of $\mathcal{F}(t)$, i. e., $\nu(t)$ directed inwards to $\mathcal{S}(t)$, the motion of the rigid body is governed by Newton’s law
(2.7) $m_{\mathcal{S}}h^{\prime\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})\nu(t)\,\mathrm{d}\Gamma,\kern 5.0pt(J\Omega)^{\prime}(t)=-\int\limits_{\partial\mathcal{S}(t)}(x-h(t))\times\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})\nu(t)\,\mathrm{d}\Gamma.$
@@
(2.8) $u^{\mathcal{F}}=0,\kern 5.0pt\partial_{\nu}Q^{\mathcal{F}}=0,\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{O},\kern 5.0pt\text{and}\kern 5.0ptu^{\mathcal{F}}=u^{\mathcal{S}},\kern 5.0pt\partial_{\nu}Q^{\mathcal{F}}=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{S}_{T}.$
@@
(2.9) $u^{\mathcal{F}}(0,x)=v_{0}(x),\kern 5.0ptQ^{\mathcal{F}}(0,x)=Q_{0}(x),\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{0},\kern 5.0pth(0)=0,\kern 5.0pth^{\prime}(0)=\ell_{0}\kern 5.0pt\text{and}\kern 5.0pt\Omega(0)=\omega_{0}.$
@@
In this paper, $\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ represents the $\mathrm{L}^{2}(\mathcal{F}_{0})$-functions with spatial average zero, while $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ denotes the solenoidal vector fields on $\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$, i. e., for the Helmholtz projection $\mathbb{P}$, we define $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\coloneqq\mathbb{P}\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$. For $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})$, $\mathrm{Y}_{\gamma}\coloneqq\mathrm{B}_{2p}^{2-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0})^{3}\cap\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathrm{B}_{2p}^{3-\nicefrac{{2}}{{p}}}(\mathcal{F}_{0},\mathbb{S}_{0}^{3})\times\mathbb{R}^{3}\times\mathbb{R}^{3}$ and $p>\nicefrac{{4}}{{3}}$, we define
(2.10) $\mathrm{X}_{\gamma}\coloneqq\{z_{0}\in\mathrm{Y}_{\gamma}:v_{0}=0,\kern 5.0pt\partial_{n}Q_{0}=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{O},\kern 5.0ptv_{0}=\ell_{0}+\omega_{0}\times y,\kern 5.0pt\partial_{n}Q_{0}=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{S}_{0}\},$
where $\mathrm{B}_{2p}^{s}(\mathcal{F}_{0})$ denotes the Besov space. The traces and normal derivatives are well-defined due to $p>\nicefrac{{4}}{{3}}$, see e. g. [44, Theorem 4.7.1]. In the main results, we will assume that $p>4$ in order to establish suitable estimates of the nonlinear terms. These estimates, carried out in detail in Section 7 and Section 8, are also based on Sobolev embeddings, which require that $p>4$. If $p\in(1,\nicefrac{{4}}{{3}})$, then we set $\mathrm{X}_{\gamma}\coloneqq\{z_{0}\in\mathrm{Y}_{\gamma}:v_{0}\cdot n=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{O},\kern 5.0ptv_{0}\cdot n=(\ell_{0}+\omega_{0}\times y)\cdot n,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{S}_{0}\}$, where $n$ denotes the outer unit normal vector to the fluid boundary, so it is directed towards $\partial\mathcal{S}_{0}$.